solve and graph 9x+6y<18
Answer by Alan3354(69443)

solve and graph 9x+6y<18
----------------------
It can be graphed, but not solved.
The graph is the area below the line, not including the line.
Graph the (straight) line by plotting 2 point of 9x+6y=18, or 3x+2y = 6, same thing.
The line has a negative slope, m = -2/3, and all of the area below the line satisfies the inequality.